# The Superclouds of the Local Milky Way

### v1

@@
11institutetext: University of Vienna, Department of Astrophysics, Türkenschanzstrasse 17, 1180 Wien, Austria 22institutetext: Istituto di Astrofisica e Planetologia Spaziali (IAPS), INAF, Via Fosso del Cavaliere 100, 00133 Roma, Italy 33institutetext: Max-Planck Institut für Astrophysik, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany 44institutetext: Ludwig-Maximilians-Universität München (LMU), Geschwister-Scholl-Platz 1, 80539 München, Germany 55institutetext: Deutsches Zentrum für Astrophysik, Postplatz 1, 02826 Görlitz, Germany 66institutetext: Excellence Cluster ORIGINS, Boltzmannstr. 2, 85748 Garching, Germany
@@
Lilly A. Kormann The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    João Alves The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Juan D. Soler The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Michelangelo Pantaleoni González The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Cameren Swiggum The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Torsten A. Enßlin The Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way    Gordian Edenhofer The Superclouds of the Local Milky WayThe Superclouds of the Local Milky Way
(Received …; accepted ..)
Recent 3D dust maps of the local Milky Way are revolutionizing our understanding of the Sun’s Galactic neighborhood, bringing much-needed insight into the large-scale organization of the interstellar medium. Focusing on the largest scales in Gaia-based 3D dust maps, we find a pattern of seven highly elongated, mostly parallel structures in the local $5\,\mathrm{kpc}^{2}$, five of which were previously unknown. These structures show pitch angles of $33.5\pm 4.0\degr$, and masses ranging from $10^{5}$ to $10^{6}$ M. We refer to these structures as superclouds. Nearly all known star-forming regions in the solar neighborhood lie within the superclouds, primarily along their central axes, supporting the idea that they are precursors to giant molecular clouds. All but one of the seven superclouds show an underlying undulation, indicating that this is not a property unique to the Radcliffe wave. We find that while the superclouds have line masses that vary by about a factor of four, their volume densities vary by about 10% only. This suggests that superclouds self-regulate their physical sizes and internal structure to maintain pressure equilibrium with their environment. These findings establish a new framework for understanding how large-scale Galactic structures shape the conditions for star formation in the solar vicinity, and likely in galaxies like the Milky Way.
@@
ISM: clouds – ISM: structure – dust, extinction
@@ 1 Introduction
The interstellar medium (ISM) of the Milky Way is a complex and structured environment characterized by features spanning a wide range of spatial scales. Among these, superclouds represent some of the most massive components. The concept of a supercloud first appeared in the seminal work of Elmegreen & Elmegreen (1983), who highlighted the regular distribution of star complexes and their associated neutral atomic hydrogen (Hi) superclouds along galactic spiral arms in nearby galaxies. They proposed that these large-scale structures, with typical masses between $10^{6}$ and $10^{7}$ M and lengths on the order of kiloparsecs, originate from gravitational or magneto-gravitational instabilities in spiral arms. This provided a link between galactic dynamics and the formation of ISM structures, motivating further investigations into the processes that govern the assembly of massive gas concentrations in galaxies.
@@
Expanding on this groundwork, Elmegreen & Elmegreen (1987) identified Hi superclouds in the inner Galaxy and noted their spatial correlation with giant molecular cloud complexes. Similarly, Dame et al. (1986), using carbon monoxide (CO) observations, revealed large molecular complexes that are embedded within superclouds. Grabelsky et al. (1987) conducted a large-scale CO survey of the Carina arm, demonstrating that molecular clouds are additional Population I tracers for this spiral arm, and established a strong link between their distribution and kinematics with those of neutral atomic hydrogen. Lada et al. (1988) presented coordinated HI and CO observations of M31 at the same spatial and spectral resolution to resolve individual giant molecular clouds (GMCs). They find that the M31 GMCs are similar to the Milky Way GMCs and embedded in a large HI spiral arm filamentary feature (see their Figure 1).
@@
Recently, Park et al. (2023) identified several Hi superclouds in the Carina arm associated with H II regions, indicating ongoing star formation within these larger atomic structures, presumably occurring within the embedded molecular clouds. Their analysis of star formation rate surface density also suggests a strong correlation with molecular gas surface density, but little correlation with atomic hydrogen surface density, highlighting the critical role of the molecular phase in the process of star birth.
@@
Subsequent theoretical work on the origin of large-scale ISM structures in galaxies can be found in Franco et al. (2002). They explored the Parker instability in the context of spiral arms, demonstrating via MHD simulations that it can lead to the formation of large-scale gas condensations, thus potentially superclouds. These structures are distributed with a characteristic wavelength along the arm and are preferentially positioned alternately above and below the galactic midplane. Kim & Ostriker (2001, 2002) further emphasized the crucial roles of self-gravity and magnetic fields as the key to gravitational runaway, and explored the formation of large-scale spiral arm features such as spurs, and their fragmentation into GMCs. There is a general consensus from observations and theory that favors a “top-down” (instability), rather than “bottom-up” (coagulation) process for the formation of at least the most massive clouds ($\gtrsim 10^{5}-10^{6}$ M), where most of the molecular material and star formation are found (Kim et al., 2003; McKee & Ostriker, 2007).
@@
Despite their recognition on galactic scales, the properties of superclouds in the immediate solar neighborhood remain poorly constrained. Several factors contribute to this knowledge gap. The isolation of large-scale complexes from the diffuse and extended background of the ISM is a challenge. Identifying individual clouds in Hi is especially difficult due to its widespread and diffuse nature. Combined with line-of-sight overlap, this complicates the establishment of clear structure boundaries. While CO serves as an effective tracer of dense molecular gas, it fails to capture the more diffuse components of the ISM, limiting our ability to trace the full extent of large-scale gaseous structures. Moreover, kinematic distance ambiguities, particularly toward the inner Galaxy, have limited the effectiveness of spectroscopic surveys (e.g., Dame et al., 1986; Elmegreen & Elmegreen, 1987). The possibility that the Sun is embedded within the remnants of an ancient supercloud, as suggested by Olano (2016), further complicates identification. Finally, the definition of a supercloud, whether it emphasizes atomic (Hi) or molecular gas (as traced by CO), varies between studies, making it difficult to establish a uniform and comprehensive survey of such structures in the local Galactic environment.
@@
The recent discovery of the Radcliffe wave, a 2.7 kiloparsec-long chain of dense gas clouds in the solar neighborhood (Alves et al., 2020), has renewed interest in the study of large-scale structures in the ISM. This remarkable structure exhibits an unexpected wavelike shape with a maximum amplitude of about 160 pc, and had remained elusive despite its size and close proximity to the Sun (within 300 pc at its closest point). Like a supercloud, the Radcliffe wave is composed of several massive star-forming regions (SFRs) embedded in well-known GMCs connected by less dense (likely atomic) gas. Its surprising oscillation about the Galactic plane and radial drift away from the Galactic Center (Konietzka et al., 2024) provides a nearby example of a large-scale Galactic-size gaseous feature, but raises questions about the uniqueness of this particular gas arrangement.
@@
To address these challenges, this paper presents a new study of superclouds in the local $\sim 5\,\mathrm{kpc}^{2}$ of the Milky Way. Our approach takes advantage of recent advances in three-dimensional dust mapping from large-scale photometric surveys, particularly those incorporating ESA Gaia data (e.g., Green et al., 2019; Lallement et al., 2019; Chen et al., 2019; Leike et al., 2020; Edenhofer et al., 2024c). By analyzing the 3D distribution of interstellar dust, an effective tracer of total gas column density (e.g., Lada et al., 1994; Alves et al., 1999; Lombardi et al., 2006), we identify regions of enhanced density and find a well-defined pattern of large-scale cloud structures. This method offers a complementary perspective to traditional kinematic analyses and avoids many of the long-standing issues, such as velocity crowding, line-of-sight confusion, and distance uncertainties.
@@
In this paper, we segment the 3D dust map of Edenhofer et al. (2024c), and construct the first catalog of superclouds within the local kiloparsec of the Galaxy. We characterize their physical properties, including filling factor, length, width, height, orientation, and mass, on the basis of the 3D dust distribution, and find a pattern of seven highly elongated, mostly parallel kiloparsec-long superclouds of the local Milky Way. We examine the connection between these superclouds and the broader ISM in the solar neighborhood, and explore potential associations with known stellar populations and other ISM components. Overall, our results provide new insights into the distribution and properties of the most massive interstellar structures in the local Galaxy, with implications for Galactic structure and the early stages of giant molecular cloud formation.
@@
Figure 1: Face-on view of the 40 HOP clouds (left) and their corresponding orientation (right). Each cloud is assigned a distinct color, the length of the line corresponds to the calculated length of the cloud. The Sun is represented by the yellow dot in the center, the Galactic Center is located to the right. An interactive 3D version of the 40 HOP clouds is available here.
@@ 2 Data
We use the posterior mean of the 3D Cartesian grid dust map interpolated by Edenhofer et al. (2024c) (hereafter E24), extending out to 1.25 kpc from the Sun. The map is based on stellar distance and extinction estimates from Zhang et al. (2023). To obtain the reconstruction, the logarithm of the dust extinction is modeled using a Gaussian process, implemented in the NIFTy.re framework (Edenhofer et al., 2024b), which incorporates smoothness priors to counteract the ”fingers-of-god” effect. This is enabled by a new generative method called Iterative Charted Refinement (ICR; Edenhofer et al., 2022). ICR efficiently models GPs on nearly arbitrarily spaced voxels, iteratively refining the map’s resolution until a target resolution is reached. The original reconstruction is performed on logarithmically spaced HEALPix spheres with decreasing angular resolution with distance from the Sun. For the posterior mean of the Cartesian interpolated map, we have a resolution of $2^{3}$ pc3, with the data given as unitless extinction values. To convert the data to hydrogen volume density, we apply
@@
where $A_{i}$ is the extinction value in the $i$-th voxel, correlated with hydrogen volume density $n_{i}$ in units of cm-3 via a factor of 1653 cm-3. The derivation of the conversion formula is detailed in O’Neill et al. (2024). By applying this conversion, we obtain a 3D hydrogen volume density distribution map.
@@ 3 Methods
To analyze and define the large-scale structures in the dust map, four different methods are tested. The Scikit-Image package (van der Walt et al., 2014), dendrograms, via the astrodendro package (Robitaille et al., 2019), HOP (Colman et al., 2024), a package developed for the extraction of molecular clouds in numerical simulations, and a deep learning approach modified at the Research Center Data Science of the University of Vienna (DS@univie). Ultimately, the HOP algorithm was selected as the best fit for this project. A discussion of the alternative methods and the reasons for their rejection can be found in Appendix A.
@@ 3.1 HOP Structure Finder
The structure finding algorithm by Colman et al. (2024) is based on the HOP clump finder of Eisenstein & Hut (1998). HOP was originally designed to identify groups in N-body simulations, and was later adapted by Colman et al. (2024) to extract and compare molecular clouds in numerical simulations. HOP depends on an externally set density threshold and peak density for the data segmentation, which makes a solely data-driven approach impossible. However, it works remarkably well given the large size of the dust map compared to the other methods, and the four segmentation parameters are well-suited for the problem.
@@ 3.1.2 Parameter Selection
To find large-scale structures in the E24 dust map, the following HOP parameters are selected after multiple test runs: density_threshold = 1.3 cm-3, density_saddle = 1.3 cm-3, density_peak = 6 cm-3, and mincells = 60000.
The density_threshold was chosen to recover and distinguish the largest present structures, while avoiding merging across the entire map due to low density connections. Increasing density_threshold does not significantly impact the identified clouds, except for slimming them. At higher values, for both density_threshold and density_peak, structures start to disappear, and only the densest regions remain. The parameter density_saddle controls whether structures are separated or merged based on the density of their connecting points. In this work, it is set equal to density_threshold to support the identification of large-scale structures. The purpose of density_saddle is to prevent the connection of structures via thin, low-density bridges, which is not an apparent problem in the final segmentation. Increasing density_saddle results in a more fragmented segmentation, breaking apart coherent clouds, and excluding otherwise present clouds of interest. Lastly, mincells is set to filter out smaller structures that might meet the other criteria, but are not of interest in this study.
Eisenstein & Hut (1998) demonstrated that the density threshold parameter has the greatest influence on the resulting group properties. Changing the initial density cut redefines the outer boundaries of the structures and determines how many voxels are included in the segmentation. In contrast, the remaining parameters primarily influence the internal structure and merging behavior.
@@ 3.2.1 Mass Calculation
For the mass calculation, we use
$M=1.37\,m_{p}\sum_{i}n_{i}\,dv_{i}$ (2)
from O’Neill et al. (2024). Here, $m_{p}$ is the proton mass, the factor 1.37 is used to convert from hydrogen mass to total mass (Heiderman et al., 2010), and $dv_{i}$ denotes the volume of each voxel. For the Cartesian map, the volume of each voxel is $2^{3}$ pc3.

### v3

@@
11institutetext: University of Vienna, Department of Astrophysics, Türkenschanzstrasse 17, 1180 Wien, Austria 22institutetext: Max-Planck Institut für Astrophysik, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany 33institutetext: Ludwig-Maximilians-Universität München (LMU), Geschwister-Scholl-Platz 1, 80539 München, Germany 44institutetext: Deutsches Zentrum für Astrophysik, Postplatz 1, 02826 Görlitz, Germany 55institutetext: Excellence Cluster ORIGINS, Boltzmannstr. 2, 85748 Garching, Germany
@@
Lilly A. Kormann    João Alves    Michelangelo Pantaleoni González
Cameren Swiggum
   Torsten A. Enßlin    Gordian Edenhofer
(Received 17 July 2025 / Accepted 3 December 2025)
Recent 3D dust maps of the local Milky Way are revolutionizing our understanding of the Sun’s Galactic neighborhood, providing much needed insight into the large-scale organization of the interstellar medium. Focusing on the largest scales in Gaia-based 3D dust maps, we find a pattern of seven highly elongated, mostly parallel structures in the local $\sim 5\,\mathrm{kpc}^{2}$, five of which were previously unknown. These structures show pitch angles of $33.5\pm 4.0\degr$ and masses ranging from $10^{5}$ to $10^{6}$ M. We refer to these structures as superclouds. Nearly all known star-forming regions in the solar neighborhood lie within the superclouds, primarily along their central axes, supporting the idea that they act as gas reservoirs for the formation of giant molecular clouds. All but one of the seven superclouds show an underlying undulation, indicating that this is not a property unique to the Radcliffe Wave. We find that while the superclouds have linear masses that vary by about a factor of 4, their volume densities only vary by about 10%. This suggests that superclouds self-regulate their physical sizes and internal structure to maintain pressure equilibrium with their environment. These findings establish a new framework for understanding how large-scale Galactic structures shape the conditions for star formation in the solar vicinity, and likely in galaxies like the Milky Way.
@@
ISM: clouds – dust, extinction – ISM: structure
@@ 1 Introduction
The interstellar medium (ISM) of the Milky Way is a complex and structured environment characterized by features spanning a wide range of spatial scales. Among these, superclouds represent some of the most massive components. The concept of a supercloud first appeared in the seminal work of elmegreen1983, who highlighted the regular distribution of star complexes and their associated neutral atomic hydrogen (Hi) superclouds along galactic spiral arms in nearby galaxies. They proposed that these large-scale structures, with typical masses between $10^{6}$ and $10^{7}$ M and lengths on the order of a few 100 pc to 1 kpc, originate from gravitational or magneto-gravitational instabilities in spiral arms. This provided a link between galactic dynamics and the formation of ISM structures, motivating further investigations into the processes that govern the assembly of massive gas concentrations in galaxies.
@@
Expanding on this groundwork, Elmegreen1987 identified Hi superclouds in the inner Galaxy and noted their spatial correlation with giant molecular cloud (GMC) complexes. Similarly, Dame1986, using carbon monoxide (CO) observations, revealed large molecular complexes that are embedded within superclouds. Grabelsky1987 conducted a large-scale CO survey of the Carina arm, demonstrating that molecular clouds are additional Population I tracers for this spiral arm, and established a strong link between their distribution and kinematics with those of neutral atomic hydrogen. Lada1988 presented coordinated Hi and CO observations of M31 at the same spatial and spectral resolution to resolve individual GMCs. They found that the M31 GMCs are similar to the Milky Way GMCs and are embedded in a large Hi spiral arm filamentary feature (see their Fig. 1).
@@
Recently, park2023 identified several Hi superclouds in the Carina arm associated with Hii regions, indicating ongoing star formation within these larger atomic structures, presumably occurring within the embedded molecular clouds. Their analysis of the star formation rate surface density also suggests a strong correlation with the molecular gas surface density, but little correlation with the atomic hydrogen surface density, highlighting the critical role of the molecular phase in the process of stellar birth.
@@
Subsequent theoretical work on the origin of large-scale ISM structures in galaxies was conducted by Franco2002. They explored the Parker instability in the context of spiral arms, demonstrating via magnetohydrodynamic simulations that it can lead to the formation of large-scale gas condensations and thus, potentially, superclouds. These structures are distributed with a characteristic wavelength along the arm and are preferentially positioned alternately above and below the galactic midplane. Kim2001; kim2002 further emphasized the roles of self-gravity and magnetic fields as the key to gravitational runaway, and explored the formation of large-scale spiral arm features such as spurs and their fragmentation into GMCs. There is a general consensus from observations and theory that favors a “top-down” (instability) rather than “bottom-up” (coagulation) process for the formation of at least the most massive clouds ($\gtrsim 10^{5}-10^{6}$ M), where most of the molecular material and star formation are found (kim2003; mckee2007).
@@
Despite their recognition on galactic scales, the properties of superclouds in the immediate solar neighborhood remain poorly constrained. Several factors contribute to this knowledge gap. The isolation of large-scale complexes from the diffuse and extended background of the ISM is a challenge. Identifying individual clouds in Hi is especially difficult due to its widespread and diffuse nature. Combined with line-of-sight overlap, this complicates the establishment of clear structure boundaries. While CO serves as an effective tracer of dense molecular gas, it fails to capture the more diffuse components of the ISM, limiting our ability to trace the full extent of large-scale gaseous structures. Moreover, kinematic distance ambiguities, particularly toward the inner Galaxy, have limited the effectiveness of spectroscopic surveys (e.g., Dame1986; Elmegreen1987). The possibility that the Sun is embedded within the remnants of an ancient supercloud, as suggested by Olano2016, further complicates identification. Finally, the definition of a supercloud, whether it emphasizes atomic (Hi) or molecular gas (as traced by CO), varies between studies, making it difficult to establish a uniform and comprehensive survey of such structures in the local Galactic environment.
@@
The recent discovery of the Radcliffe Wave, a 2.7 kiloparsec-long chain of dense gas clouds in the solar neighborhood (alves2020), has renewed interest in the study of large-scale structures in the ISM. This remarkable structure exhibits an unexpected wavelike shape with a maximum amplitude of about 160 pc and had remained elusive despite its size and close proximity to the Sun (within 300 pc at its closest point). Like a supercloud, the Radcliffe Wave is composed of several massive star-forming regions (SFRs) embedded in well-known GMCs and connected by less dense (likely atomic) gas. Its surprising oscillation about the Galactic plane and radial drift away from the Galactic Center (konietzka2024) provides a nearby example of a large-scale Galactic-size gaseous feature but raises questions about the uniqueness of this particular gas arrangement.
@@
To address these challenges, we present a study of newly identified superclouds in the local $\sim 5\,\mathrm{kpc}^{2}$ of the Milky Way. Our approach takes advantage of recent advances in 3D dust mapping from large-scale photometric surveys, particularly those incorporating ESA Gaia data (e.g., Green2019; Lallement2019; Chen2019; Leike2020; Edenhofer2024). By analyzing the 3D distribution of interstellar dust, an effective tracer of the total gas column density (e.g., lada1994; alves1999; lombardi2006), we identified regions of enhanced density and find a well-defined pattern of large-scale cloud structures. This method offers a complementary perspective to traditional kinematic analyses and avoids many of the longstanding issues, such as velocity crowding, line-of-sight confusion, and distance uncertainties.
@@
In this study we segmented the 3D dust map of Edenhofer2024 and constructed the first catalog of superclouds within the local kiloparsec of the Galaxy. We characterized their physical properties, including filling factor, length, width, height, orientation, and mass, on the basis of the 3D dust distribution, and find a pattern of seven highly elongated, mostly parallel kiloparsec-long superclouds in the local Milky Way. We examined the connection between these superclouds and the broader ISM in the solar neighborhood and explored potential associations with known stellar populations and other ISM components. Overall, our results provide new insights into the distribution and properties of the most massive interstellar structures in the local Galaxy, with implications for Galactic structure and the early stages of GMC formation.
@@
Figure 1: Face-on view of the 40 HOP clouds (left) and their corresponding orientation (right). Each cloud is assigned a distinct color, and the length of the line corresponds to the calculated length of the cloud. The Sun is represented by the yellow dot in the center; the Galactic Center is located to the right. An interactive 3D version of the 40 HOP clouds is available here.
@@ 2 Data
We used the posterior mean of the 3D Cartesian grid dust map interpolated by Edenhofer2024, which extends out to 1.25 kpc from the Sun. The map is based on stellar distance and extinction estimates from zhang2023. To obtain the reconstruction, the logarithm of the dust extinction is modeled using a Gaussian process, implemented in the NIFTy.re framework (edenhofer2024_nifty), which incorporates smoothness priors to counteract the ”fingers-of-god” effect. This is enabled by a new generative method called iterative charted refinement (edenhofer2022_icr), which efficiently models Gaussian processes on nearly arbitrarily spaced voxels, iteratively refining the map’s resolution until a target resolution is reached. The original reconstruction is performed on logarithmically spaced HEALPix spheres with decreasing angular resolution with distance from the Sun.
For the posterior mean of the Cartesian interpolated map, we have a resolution of $2^{3}$ pc3, with the data given as unitless extinction values. To convert the data to hydrogen volume density, we applied
@@
where $A_{i}$ is the extinction value in the $i$-th voxel, correlated with hydrogen volume density $n_{i}$ in units of cm-3 via a factor of 1653 cm-3. The derivation of the conversion formula is detailed in oneill2024. By applying this conversion, we obtained a 3D hydrogen volume density distribution map.
@@ 3 Methods
To analyze and define the large-scale structures in the dust map, four different methods are tested: the scikit-image package (scikitimage); dendrograms via the astrodendro package (astrodendro); HOP (colman2024), a package developed for the extraction of molecular clouds in numerical simulations; and a deep learning approach modified at the Research Center Data Science of the University of Vienna (DS/Univie). Ultimately, the HOP algorithm was selected as the best fit for this project. A discussion of the alternative methods and the reasons for their rejection can be found in Appendix A.
@@ 3.1 HOP structure finder
The structure finding algorithm by colman2024 is based on the HOP clump finder of eisenstein1998. HOP was originally designed to identify groups in N-body simulations, and was later adapted by colman2024 to extract and compare molecular clouds in numerical simulations.
HOP depends on an externally set density threshold and peak density for the data segmentation, which precludes a solely data-driven approach. However, it works remarkably well given the large size of the dust map compared to the other methods, and the four segmentation parameters are well suited for the problem.
@@ 3.1.2 Parameter selection
To find large-scale structures in the Edenhofer2024 dust map, the following HOP parameters were selected after multiple test runs: density_threshold = 1.3 cm-3, density_saddle = 1.3 cm-3, density_peak = 6 cm-3, and mincells = 60.000.
The density_threshold was chosen to recover and distinguish the largest present structures, while avoiding merging across the entire map due to low-density connections. Multiple parameter tests with density_threshold values between 1 cm-3 and 5 cm-3 have shown that the exact value of the threshold does not significantly impact the general trend of the recovered clouds. While the threshold itself is below the resolution limit at which individual low-density structures can be reliably resolved, the integrated extinction measurements constrain the dust distribution well enough to identify large-scale structures with lower densities, even if their boundaries cannot be precisely traced on a parsec scale. At higher values, for both density_threshold and density_peak, structures start to disappear, and only the densest regions remain. The parameter density_saddle controls whether structures are separated or merged based on the density of their connecting points. In this work, it was set equal to density_threshold to support the identification of large-scale structures. The purpose of density_saddle is to prevent the connection of structures via thin, low-density bridges, which is not an apparent problem in the final segmentation. Increasing density_saddle results in a more fragmented segmentation, breaking apart coherent clouds, and excluding otherwise present clouds of interest. Lastly, mincells is set to filter out smaller structures that might meet the other criteria, but are not of interest in this study.
eisenstein1998 demonstrated that the density threshold parameter has the greatest influence on the resulting group properties. Changing the initial density cut redefines the outer boundaries of the structures and determines how many voxels are included in the segmentation. In contrast, the remaining parameters primarily influence the internal structure and merging behavior.
@@ 3.2.1 Mass calculation
For the mass calculation, we used
$M=1.37\,m_{p}\sum_{i}n_{i}\,\text{d}v_{i}$ (2)
from oneill2024. Here, $m_{p}$ is the proton mass, the factor 1.37 is used to convert from hydrogen mass to total mass (heiderman2010), and $\text{d}v_{i}$ denotes the volume of each voxel. For the Cartesian map, the volume of each voxel is $2^{3}$ pc3.